### v1

# Learning Geometric-Aware Quadrature Rules for Functional Minimization

Abstract.
Accurate numerical integration over non-uniform point clouds is a challenge for modern mesh-free machine learning solvers for partial differential equations (PDEs) using variational principles. While standard Monte Carlo (MC) methods are not capable of handling a non-uniform point cloud, modern neural network architectures can deal with permutation-invariant inputs, creating quadrature rules for any point cloud. In this work, we introduce QuadrANN, a Graph Neural Network (GNN) architecture designed to learn optimal quadrature weights directly from the underlying geometry of point clouds. The design of the model exploits a deep message-passing scheme where the initial layer encodes rich local geometric features from absolute and relative positions as well as an explicit local density measure. In contrast, the following layers incorporate a global context vector. These architectural choices allow the QuadrANN to generate a data-driven quadrature rule that is permutation-invariant and adaptive to both local point density and the overall domain shape. We test our methodology on a series of challenging test cases, including integration on convex and non-convex domains and estimating the solution of the Heat and Fokker-Planck equations. Across all the tests, QuadrANN reduces the variance of the integral estimation compared to standard Quasi-Monte Carlo methods by warping the point clouds to be more dense in critical areas where the integrands present certain singularities. This enhanced stability in critical areas of the domain at hand is critical for the optimization of energy functionals, leading to improved deep learning-based variational solvers.
Key words and phrases:
Numerical integration, Quadrature rules, Graph Neural Networks, Deep learning, Partial differential equations, Variational methods, Mesh-free numerical methods
2020 Mathematics Subject Classification:
65D32,65N12,68T07
1. Introduction
Variational principles are fundamental to the mathematics and physical sciences, creating the ground for solving challenging problems in fields covering classical and statistical mechanics to modern probability theory. These principles often require the minimization a functional, which typically involves an integral over a particular domain $\Omega$:
$\mathcal{J}[u]=\int_{\Omega}L(x,u(x),\nabla u(x))dx,\quad u\in H^{1}(\Omega),$ (1)
where $L$ is a given operator. The objective is to find a function $u^{*}\in H^{1}(\Omega)$ that minimizes this functional, i.e., $\mathcal{J}[u^{*}]\leq\mathcal{J}[u]$ for any function $u\in H^{1}(\Omega)$. Such a problem is equivalent to solving the corresponding Euler-Lagrange partial differential equation (PDE), and the pursuit of its solution is a cornerstone of computational science.
For many years, the numerical solution of these problems has been driven by mesh-based methods like the Finite Element Method (FEM)[1]. While powerful these methods are, they heavily rely on domain meshing, which can be computationally restrictive for problems in high-dimensional spaces or with complex geometries. In recent years, we have witnessed a shift in the methodologies that are used to solve problems, thanks to the introduction of deep learning-based solvers. Methods spanning the Deep Ritz Method (DRM) [2] and Physics-Informed Neural Networks (PINNs) [3] have been introduced as mesh-free alternatives. These modern approaches reformulate the variational problem as an optimisation problem, where now the parameters of an Artificial Neural Network (ANN), representing the solution $u$, are trained to minimize the functional $\mathcal{J}[u]$.
A significant aspect of deep learning solvers is the accurate numerical approximation of the integral shown in Equation (1). The solving process depends heavily on the minimization of the discretized functional, $\hat{\mathcal{J}}[u]$, typically evaluated with respect to a given set of sample points. The accuracy of this numerical integration is critical for obtaining an accurate solution. As this work will underscore, unless the approximation $\hat{\mathcal{J}}[u]$ is highly accurate, the minimizer $\hat{u}^{*}$ obtained from training the ANN will generally not be a reasonable estimate of the actual minimizer $u^{*}$ of the original functional.
The simplest and common approach for approximating such integrals is, of course, the Monte Carlo (MC) method. In an MC-driven approach, the integral is estimated by averaging the values of the integrand over a set of randomly sampled points. While the universal MC approach is this method suffers from slow convergence, with an error rate of $\mathcal{O}(N^{-1/2})$, requiring a large number of sample points $n$ for acceptable accuracy. Variants like Quasi-Monte Carlo (QMC) methods offer faster convergence for certain classes of integrands [4]. However, QMC methods may not be optimal for the complex/non-linear integrands that often arise when training neural networks to solve high-dimensional PDEs, a challenge this work aims to address.
This challenge has motivated research into more sophisticated quadrature strategies. For instance, Bayesian Quadrature models the integrand as a Gaussian process to provide a posterior distribution over the value of the integral, which can be more sample-efficient [5]. Additionally, the concept of adaptive quadrature that concentrates more points to areas where the integrand is complex or varies rapidly, is well-established in classical numerical analysis [6] and is beginning to find its way into deep learning contexts.
In this work, we claim that for variational problems solved with Artificial Neural Networks (ANNs), a static quadrature rule is generally insufficient. Modern deep learning solvers often employ non-uniform point distributions that are chosen adaptively to minimize the PDE residual. This dynamic nature of point sampling necessitates a quadrature rule that can also adapt to arbitrary non-uniformity.
We therefore develop a ANN-based approach for numerical integration, focusing on learning optimal quadrature rules directly from the sample points. Instead of relying on predefined static weights, we propose a model that learns a direct mapping from a set of points to a corresponding set of optimal quadrature weights. This is achieved by training the ANN to produce weights that guarantee the exact (or almost exact) integration of a set of basis functions, creating a highly accurate and data-driven quadrature rule.
Our work builds upon a growing interest in machine learning for numerical integration. This field includes probabilistic methods like Bayesian Quadrature, which models the integrand as a Gaussian process to estimate the uncertainty [7] of the integral, and direct approaches that train neural networks to predict optimal quadrature weights for specific function classes or domains [2, 8, 9]. Unlike methods that focus on learning adaptive sampling policies, our approach learns a direct mapping from a point cloud configuration to its corresponding weights. Our method is particularly suited for the non-uniform sampling policies inherent in modern deep learning solvers that aim to close the gap between the theoretical functional and its practical, discretised approximations.
The remainder of this work is as follows. In Section 2, we recap the neural quadrature problem for functional integrals. Section 3 introduces our approach for generating non-uniform point clouds model for model training and testing. In Section 4, we present the proposed QuadrANN architecture and analyze our architectural decisions. Section 5 presents applications to demonstrate the performance of QuadrANN, and finally, Section 6 provides the conclusion.
2. Neural Quadrature for Functional Integrals
Let domain $\Omega\subset\mathbb{R}^{d}$. The core idea in a quadrature rule is to estimate an integral using a weighted sum over a set of $n$ points $\mathcal{X}=\{x_{1},x_{2},\dots,x_{n}\}$ sampled from $\Omega$:
$\int_{\Omega}L(x,u(x),\nabla u(x))dx\approx\sum_{i=1}^{N}w_{i}L(x_{i},u(x_{i}),\nabla u(x_{i})).$ (2)
The main challenge here is to find the optimal weights $(w_{1},\dots,w_{n})$ for a given set of points $\mathcal{X}$. These weights should depend on the overall geometry of the point cloud but be independent of the ordering in which the points are provided to a model. To achieve this, we represent the mapping from the point set $\mathcal{X}$ to the weights $(w_{i})_{i=1}^{n}$ by means of an Artificial Neural Network (ANN) that is insensitive to the permutation of its inputs. We denote this network as $W(\cdot\ ;\theta)$, which takes the entire set $\mathcal{X}$ as input and produces $N$ corresponding weights. The integral approximation is then:
$\int_{\Omega}L(x,u(x),\nabla u(x))dx\approx\sum_{i=1}^{n}W_{i}(\mathcal{X};\theta)L(x_{i},u(x_{i}),\nabla u(x_{i})),$ (3)
where $W_{i}(\cdot\ ;\theta)$ is the $i$-th output of the network corresponding to the point $x_{i}$. While the underlying framework is applicable to more general domains, the theoretical development of our ANN quadrature rule will henceforth be mainly focused on the $d$-dimensional hypercube $\Omega=[0,1]^{d}$. However, we will expand our analysis to include more complex geometries, such as a non-convex L-shaped domain in 2D (see Subsection 5.1.2).
The network is trained to produce model parameters that guarantee a good integration performance for a predefined basis set of test functions. This basis is designed to be highly versatile, combining polynomial and trigonometric functions to effectively approximate a broad class of integrands. The core of the basis consists of multivariate polynomials of total degree at most $J$, which is fixed at $5$ throughout this work, formed by tensor products of normalized Hermite polynomials, $\{\Phi_{\bm{j}}(\cdot)\}$:
$\Phi_{\bm{j}}(x)=\prod_{k=1}^{d}\tilde{H}_{e_{j_{k}}}(x_{k}),\ \text{where}\sum_{k=1}^{d}j_{k}\leq J.$ (4)
The polynomials $\tilde{H}_{e_{j_{k}}}$ are normalised such that their integral over $[0,1]$ is unity. To increase the capability of the basis for oscillatory functions, the degree-zero Hermite polynomial is reformulated as a trigonometric term:
$\tilde{H}_{e_{0}}(x;\kappa,\varphi)=1+\cos(\kappa\pi x+\varphi)-\int_{\Omega}\cos(\kappa\pi x+\varphi)dx,$ (5)
where $\kappa$ is an integer randomly sampled from $\{1,2,\dots,K_{\mathrm{max}}\}$ during training and $\varphi\in[0,2\pi)$ is a random phase . This hybrid construction, controlled by hyperparameters $J$ and $K_{\mathrm{max}}$, creates a powerful approximation space. The polynomial components are well-suited for capturing smooth, low-frequency behaviour, while the randomized trigonometric terms are explicitly included to represent oscillatory patterns that are difficult to approximate efficiently with a pure polynomial basis. It is important to emphasize that this reformulation does not affect the ability of the model to integrate constants. This property is instead enforced by the architecture of the network, as the final softmax activation layer constrains the output weights to sum to unity ($\sum W_{i}=1$).
The training objective is operationalized through the following discrete loss function, which penalizes deviation from exactness on the test basis:
$\mathcal{C}(\mathcal{X};\theta)=\sum_{\bm{j}:\sum_{k=1}^{d}j_{k}\leq J}\left(\left(\sum_{i=1}^{n}W_{i}(\mathcal{X};\theta)\Phi_{\bm{j}}(x_{i})\right)-1\right)^{2}.$ (6)
The training phase consists of multiple iterations where a new set of points $\mathcal{X}$ is obtained by resampling, and the model parameters $\theta$ are updated using the ADAM optimizer.
3. Generating Non-Uniform Point Clouds
To ensure robust training and evaluation of our neural network-based quadrature model, it is essential to be capable of generating point clouds that not only cover the entire domain but also feature non-uniform densities. This approach simulates the conditions commonly encountered in practical applications, such as adaptive mesh refinement or scenarios where physical phenomena are concentrated in specific areas. Our procedure for generating these point clouds is a two-step process that combines the superior uniformity of Quasi-Monte Carlo sequences with a deterministic, non-linear transformation.
3.1. Quasi-Random Base Sampling
Instead of using entirely random numbers, which can lead to the formation of gaps, we are generating sets of points with low-discrepancy characteristics using a Quasi-Monte Carlo (QMC) method. QMC sequences are designed to cover the sampling space more evenly than entirely random sampling strategies. Specifically, in this work, we employ the Sobol sequence[10], a widely-used standard for QMC integration.
Given $n$, we generate a base point set $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\dots,\mathbf{s}_{n}\}$, where each point $\mathbf{s}_{i}$ is drawn from the Sobol sequence within the unit hypercube $\Omega=[0,1]^{d}$. This set $\mathcal{S}$ provides a basis for our sampling method.
3.2. Non-Linear Warping
While the Sobol sequence provides a well-behaved uniform distribution, our goal is to create point clouds with varying densities. To achieve this in a controllable and general manner, we can apply a bijective, non-linear transformation $\mathbf{G}:[0,1]^{d}\to[0,1]^{d}$ to uniformly distributed points. This vector-valued function warps the entire unit hypercube, allowing for complex, non-separable distortions beyond simple coordinate-wise stretching.
This transformation deterministically modifies the local density of points. Regions where the transformation compresses space lead to higher point density, while regions that are expanded result in lower density.
Each point $\mathbf{s}_{i}\in\mathcal{S}$ is to be mapped to a transformed point $\mathbf{x}_{i}$ to create the final non-uniform point cloud $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}$. The transformation is defined as follows:
$\mathbf{x}_{i}=\mathbf{G}(\mathbf{s}_{i}),\quad\text{for }i=1,\dots,n.$ (7)
This two-step procedure enables us to create on demand non-uniform point clouds with certain characteristics by simply modifying the warping function $\mathbf{G}(\cdot)$. The resulting unordered set $\mathcal{X}$ serves then as the direct input for our quadrature model, creating a robust framework to learn quadrature weights for arbitrary point clouds.
4. The QuadrANN Architecture
To address the challenge of learning quadrature rules directly from point clouds, we introduce a new neural network architecture, QuadrANN. It is a deep Graph Neural Network (GNN) architecture [11, 12] that builds on principles from graph models for point clouds [13] and is designed to capture and utilize geometric information at multiple scales effectively. Our approach is based on the principle that the quadrature weight for a given point depends not only on its position but also on its role within the geometry and structure of the domain. The architecture is composed of three primary stages: an initial geometric encoding layer, a number of globally-aware propagation layers, and a final weight prediction neural network.
Several critical architectural decisions were made to optimize QuadrANN for this task. First, to capture rich local information, we combine two powerful features: Positional Encoding (PE), a technique popularized by Transformers and coordinate-based networks [14, 15], provides a high-frequency understanding of relative positions, while an explicit density estimate provides a direct, robust measure of local point concentration. Second, to ensure the entire shape informs the learned weights, each message-passing step is conditioned on a global context vector, a strategy used in seminal point cloud models [16]. Third, inspired by DenseNets [17], we employ a dense concatenation of feature maps from all layers. This gives the prediction network simultaneous access to multi-scale features, mitigating the over-smoothing problem, common in deep GNNs [18].
Ultimately, the complete model network is trained using a global softmax function in the final layer. This activation function guarantees that the output weights are positive and sum to one, which satisfies the mathematical properties of a quadrature rule. The following sections will provide a detailed explanation of each component of this architecture.
The complete formulation of the model is presented in Algorithm 1.
4.1. Input Feature Engineering
In order to prepare the raw point cloud, we perform an initial feature engineering on the input coordinates. The raw coordinates $\{\mathbf{x}_{i}\}_{i=1}^{n}$, assumed to lie within the unit hypercube $0,1]^{d}$ (or in a subset of this domain). We first normalize the raw coordinates to the typical domain $1,1]^{d}$ via the linear transformation $\mathbf{x}_{i}\leftarrow 2\mathbf{x}_{i}-1$.
To enable the network to learn high-frequency patterns and to improve its ability to reason about both absolute and relative positions of the points, we employ Positional Encoding (PE). Using PE is a common technique that maps a low-dimensional continuous space into a higher-dimensional feature space by employing a set of sinusoidal functions with various frequencies.
For a general input vector $\mathbf{x}\in\mathbb{R}^{d}$, the positional encoding $\text{PE}(\mathbf{x})$ is constructed by applying the encoding to each component of the vector and concatenating the results:
$\text{PE}_{p}(\mathbf{x})=(\gamma(x_{1};p),\gamma(x_{2};p),\dots,\gamma(x_{d};p))$ (8)
The encoding function $\gamma(\cdot;p)$ for a single scalar component $\alpha$ is defined as:
$\gamma(\alpha;p)=\left(\sin(2^{0}\pi\alpha),\cos(2^{0}\pi\alpha),\dots,\sin(2^{p-1}\pi\alpha),\cos(2^{p-1}\pi\alpha)\right)$
where $p$ is a hyperparameter representing the number of frequency bands. This multi-scale representation provides the model with a rich, unambiguous signal for both fine-grained detail and global position, resulting in a feature vector of dimension $d\times 2p$.
4.2. Geometric Encoding Layer
The first layer of our network creates a high-dimensional latent representation for each point $\mathbf{x}_{i}\in\mathbb{R}^{d}$. The dimensionality of these representations is denoted by $Q_{1}$, which we constrain to be an even number. An even number ensures that the feature dimension can be cleanly halved, as the final node features are formed by concatenating two vectors of equal size ($Q_{1}/2$). The resulting latent vector is denoted as $\mathbf{o}_{i}^{(1)}\in\mathbb{R}^{Q_{1}}$.
We begin by defining two $k$-NN graphs. First, a graph with $k^{\prime}$ neighbours is used to compute a local density feature, $\rho_{i}\in\mathbb{R}$, for each point. The density feature is defined to be the reciprocal of the mean distance of a point’s neighbours to their local centroid, providing a robust measure of point concentration. Specifically, for each point $\mathbf{x}_{i}$, we define its neighborhood centroid as $\mathbf{c}_{i}=\frac{1}{k^{\prime}}\sum_{j:\mathbf{x}_{j}\in V_{k^{\prime}}(\mathbf{x}_{i})}\mathbf{x}_{j}$, and the density is then $\rho_{i}=\left(\frac{1}{k^{\prime}}\sum_{j:\mathbf{x}_{j}\in V_{k^{\prime}}(\mathbf{x}_{i})}\|\mathbf{x}_{j}-\mathbf{c}_{i}\|_{2}\right)^{-1}$. For computational efficiency, we can always set $k^{\prime}=k$, requiring the $k$-NN search to be performed only once for each point $\mathbf{x}_{i}$.
By incorporating the explicit density feature, $\rho_{j}$, into the message-passing scheme, the model acquires essential information about local point concentration. Although a GNN can implicitly deduce density from its inputs, supplying a precise measure enables the model to reason about point density directly and frees up model capacity, enabling the subsequent MLP to learn the more complex relationships between geometry, density, and the resulting quadrature weights. Ultimately, this leads to a more robust and accurate quadrature rule.
Subsequently, a graph with $k$ neighbours determines the neighbourhood $V_{k}(\mathbf{x}_{i})$, which facilitates the main message passing among the points. Within this neighbourhood, a message vector, $\mathcal{G}_{ij}$, is generated by augmenting the geometric features with the pre-computed density feature of the neighbours:
$\mathcal{G}_{ij}=\left[\mathbf{x}_{j},\text{PE}_{p_{1}}(\mathbf{x}_{j}),\mathbf{x}_{j}-\mathbf{x}_{i},\text{PE}_{p_{2}}(\mathbf{x}_{j}-\mathbf{x}_{i}),\rho_{j}\right]$ (9)
This message combines five key features:
  1. (1)
    the absolute position of the neighbor ($\mathbf{x}_{j}$) and its positional encoding ($\text{PE}_{p_{1}}(\mathbf{x}_{j})$) to provide global context;
  2. (2)
    the relative position vector ($\mathbf{x}_{j}-\mathbf{x}_{i}$) to capture their Euclidean relationship;
  3. (3)
    the positional encoding of the relative vector ($\text{PE}_{p_{2}}(\mathbf{x}_{j}-\mathbf{x}_{i})$) to provide a high-frequency representation of the local relationship;
  4. (4)
    the local density explicit feature ($\rho_{j}$) at the neighboring point to provide explicit information about the local concentration of the points.
A dedicated MLP then processes this augmented data, denoted $\text{MLP}_{1}$. This network consists of three layers and its role is to transform the rich input data into a meaningful latent feature space:
  1. (1)
    a linear transformation that maps the input $\mathcal{G}_{ij}$ to a hidden representation, followed by a GELU activation function;
  2. (2)
    a second linear transformation and GELU activation further process the features;
  3. (3)
    a final linear transformation that produces the output message $\mathbf{m}_{ij}^{(1)}\in\mathbb{R}^{Q_{1}/2}$.
Finally, all learned messages are aggregated using a higher-order statistical function to form the initial node feature $\mathbf{o}_{i}^{(1)}\in\mathbb{R}^{Q_{1}}$:
$\mathbf{o}_{i}^{(1)}=\left[\bm{\mu}_{i}(\mathbf{m}_{ij}^{(1)}),\bm{\sigma}_{i}(\mathbf{m}_{ij}^{(1)})\right]$ (10)
This initial encoding provides a rich, geometrically-aware basis for subsequent layers.
4.3. Globally-Aware Propagation Layers
Following the initial encoding, a stack of L1L-1 propagation layers refines the node features. At each layer $l\in\{2,\dots,L\}$, a global context vector $\mathbf{g}^{(l-1)}\in\mathbb{R}^{Q_{1}}$ is first computed by averaging all node features from the previous layer. A message is then computed by a layer-specific $\text{MLP}_{l}$, which has a compact two-layer structure sufficient for operating on already-learned features:
  1. (1)
    a linear tranformation maps the concatenated input vector $\mathbf{o}_{j}^{(l-1)},\mathbf{o}_{j}^{(l-1)}-\mathbf{o}_{i}^{(l-1)},\mathbf{g}^{(l-1)}]$ of dimension $3Q_{1}$ to a hidden representation, followed by a GELU activation function;
  2. (2)
    a final linear transformation produces the output message $\mathbf{m}_{ij}^{(l)}\in\mathbb{R}^{Q_{1}/2}$.
The new node features, $\mathbf{o}_{i}^{(l)}\in\mathbb{R}^{Q_{1}}$, are then formed by concatenating the mean and standard deviation of the messages. After the final propagation layer, we create a comprehensive representation, $\mathbf{o}_{i}^{\text{final}}\in\mathbb{R}^{L\cdot Q_{1}}$, by concatenating the feature vectors from all layers:
$\mathbf{o}_{i}^{\text{final}}=\left[\mathbf{o}_{i}^{(1)},\mathbf{o}_{i}^{(2)},\dots,\mathbf{o}_{i}^{(L)}\right].$ (11)
4.4. Weight Prediction Network
The final stage of the model maps the comprehensive feature vector $\mathbf{o}_{i}^{\text{final}}$ to a quadrature weight $w_{i}$. This is performed by a point-wise MLP, $\text{MLP}_{\text{out}}$, whose structure is defined by a second hyperparameter, $Q_{2}$. We require again $Q_{2}$ to be an even positive integer. This network is designed as follows:
  1. (1)
    an input linear transformation takes the $L\times Q_{1}$-dimensional vector $\mathbf{o}_{i}^{\text{final}}$;
  2. (2)
    a first hidden linear layer maps the input to a $Q_{2}$-dimensional space, followed by a GELU activation and a dropout (=0.5=0.5) layer for regularization;
  3. (3)
    a second hidden layer reduces the dimension from $Q_{2}$ to $Q_{2}/2$, followed again by GELU and dropout (=0.5=0.5);
  4. (4)
    an output layer maps the final hidden representation to a single scalar logit, $v_{i}$.
The vector of all logits, $\mathbf{v}=(v_{1},\dots,v_{n})^{T}$, is then normalized using a global softmax function to ensure the weights are positive and sum to unity ($w_{i}>0,\sum_{i}w_{i}=1$).
Algorithm 1 The QuadrANN Architecture
1:Input: Point cloud $\mathcal{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}\subseteq\mathbb{R}^{d}$, density neighbors $k^{\prime}$, message neighbors $k$, layers $L$, PE bands $p_{1},p_{2}$, node dimension $Q_{1}$, prediction network parameter $Q_{2}$.
2:Learnable Parameters: Geometric MLP $\text{MLP}_{1}$, Propagation MLPs $\{\text{MLP}_{l}\}_{l=2}^{L}$, Prediction MLP $\text{MLP}_{\text{out}}$.
3:Output: Quadrature weights $\mathbf{w}\in\mathbb{R}^{n}$.
4:procedure QuadrANN($\{\mathbf{x}_{i}\}_{i=1}^{n}$)
5:  $\triangleright$ Stage 1: Geometric Encoding Layer
6:  Find density neighborhood $V_{k^{\prime}}(\mathbf{x}_{i})$ for each point $\mathbf{x}_{i}$.
7:  for each point $\mathbf{x}_{i},\ i=1,\dots,n$ do $\triangleright$ Pre-compute density feature
8:   $\mathbf{c}_{i}\leftarrow\frac{1}{k^{\prime}}\sum_{j:\mathbf{x}_{j}\in V_{k^{\prime}}(\mathbf{x}_{i})}\mathbf{x}_{j}$ $\triangleright$ Neighborhood centroid
9:   $\rho_{i}\leftarrow(\frac{1}{k^{\prime}}\sum_{j:\mathbf{x}_{j}\in V_{k^{\prime}}(\mathbf{x}_{i})}\|\mathbf{x}_{j}-\mathbf{c}_{i}\|_{2})^{-1}$ $\triangleright$ Density feature is the inverse of mean distance
10:  end for
11:
12:  Find message passing neighborhood $V_{k}(\mathbf{x}_{i})$ for each point $\mathbf{x}_{i}$. $\triangleright$ Can be skipped if $k^{\prime}=k$
13:  for each point $\mathbf{x}_{i},\ i=1,\dots,n$ do
14:   for each neighbor $\mathbf{x}_{j}\in V_{k}(\mathbf{x}_{i})$ do
15:     $\mathcal{G}_{ij}\leftarrow\left[\mathbf{x}_{j},\text{PE}_{p_{1}}(\mathbf{x}_{j}),\mathbf{x}_{j}-\mathbf{x}_{i},\text{PE}_{p_{2}}(\mathbf{x}_{j}-\mathbf{x}_{i}),\rho_{j}\right]$
16:     $\mathbf{m}_{ij}^{(1)}\leftarrow\text{MLP}_{1}(\mathcal{G}_{ij})$
17:   end for
18:   $\bm{\mu}_{i}^{(1)}\leftarrow\underset{j:\mathbf{x}_{j}\in V_{k}(\mathbf{x}_{i})}{\text{Mean}}(\{\mathbf{m}_{ij}^{(1)}\})$;  $\bm{\sigma}_{i}^{(1)}\leftarrow\underset{j:\mathbf{x}_{j}\in V_{k}(\mathbf{x}_{i})}{\text{Std}}(\{\mathbf{m}_{ij}^{(1)}\})$
19:   $\mathbf{o}_{i}^{(1)}\leftarrow[\bm{\mu}_{i}^{(1)},\bm{\sigma}_{i}^{(1)}]$
20:  end for
21:
22:  $\triangleright$ Stage 2: Globally-Aware Propagation Layers
23:  for $l=2,\dots,L$ do
24:   $\mathbf{g}^{(l-1)}\leftarrow\frac{1}{n}\sum_{p=1}^{n}\mathbf{o}_{p}^{(l-1)}$
25:   for each $i=1,\dots,n$ do
26:     for each $j:\mathbf{x}_{j}\in V_{k}(\mathbf{x}_{i})$ do
27:      $\mathbf{m}_{ij}^{(l)}\leftarrow\text{MLP}_{l}([\mathbf{o}_{j}^{(l-1)},\mathbf{o}_{j}^{(l-1)}-\mathbf{o}_{i}^{(l-1)},\mathbf{g}^{(l-1)}])$
28:     end for
29:     $\mathbf{o}_{i}^{(l)}\leftarrow[\underset{j:\mathbf{x}_{j}\in V_{k}(\mathbf{x}_{i})}{\text{Mean}}(\{\mathbf{m}_{ij}^{(l)}\}),\underset{j:\mathbf{x}_{j}\in V_{k}(\mathbf{x}_{i})}{\text{Std}}(\{\mathbf{m}_{ij}^{(l)}\})]$
30:   end for
31:  end for
32:
33:  $\triangleright$ Stage 3: Final Representation and Weight Prediction
34:  for each $i=1,\dots,n$ do
35:   $\mathbf{o}_{i}^{\text{final}}\leftarrow[\mathbf{o}_{i}^{(1)},\mathbf{o}_{i}^{(2)},\dots,\mathbf{o}_{i}^{(L)}]$
36:   $v_{i}\leftarrow\text{MLP}_{\text{out}}(\mathbf{o}_{i}^{\text{final}})$
37:  end for
38:  $\mathbf{w}\leftarrow\text{Softmax}((v_{1},\dots,v_{n})^{T})$
39:
40:  return $\mathbf{w}$
41:end procedure
5. Applications
In this section, we demonstrate the robustness of the QuadrANN framework through a number of numerical experiments. We begin by directly evaluating its performance on pure integration problems and then showcase its utility as a component within complex variational solvers for time-dependent Partial Differential Equations (PDEs).
A challenging point in numerical integration, particularly for functions with sharp gradients or localised peaks, is the fact that standard sampling methods may not adequately cover the regions of interest. Methods like Monte Carlo or even Quasi-Monte Carlo (QMC) that typically distribute points uniformly may undersample those critical parts of the domain where the contribution of the integrand is most significant, leading to slow convergence or high variance approximations. To achieve high accuracy with a limited number of points, the sampling distribution should ideally be concentrated in these important regions. Our approach is designed to excel in precisely this scenario by learning the correct quadrature weights for a given, potentially non-uniform, point distribution.
In order to ensure the robustness of our model, we test it in the context of non-uniform point clouds as described in the previous sections. These clouds are generated by first creating a low-discrepancy point set using a scrambled Sobol sequence and then applying a non-linear, coordinate-wise warping function, $\mathbf{G}(\mathbf{x})=(g(x_{1}),\dots,g(x_{d}))$. The function $g:[0,1]\to[0,1]$ is defined as:
$g(s)=0.95s+0.05\left(4(s-0.5)^{3}+0.5\right).$ (12)
This transformation intentionally concentrates points around the center of the domain, simulating a scenario where the energy of the integrand is localized around that point. By training QuadrANN on a constantly changing set of warped point clouds, the network learns a quadrature rule that is not tied to any single cloud but is instead a general-purpose function adaptable to local point density variations. Note that this exact transformation is used for all examples that follow.
Our numerical examples are divided into two parts. First, we present numerical integration examples to directly benchmark the accuracy and stability of QuadrANN against standard Sobol-QMC. These tests include a simple 2D unit square, a non-convex 2D L-shaped domain, and a 4D hypercube to test the performance of the model under the curse of dimensionality. Second, we apply QuadrANN to solve PDEs: the Heat equation in 2D and the Fokker-Planck equation in 4D. These examples demonstrate the practical advantage of using an adaptive, data-driven quadrature rule within variational, time-stepping schemes that are common in modern scientific computing.
5.1. Numerical Integration
To test the performance of QuadrANN, we conducted a number of numerical examples. For each example, we trained a dedicated model to learn a quadrature rule for a specific domain and according to a certain point cloud generation policy. In particular, at each training epoch, we generate a batch of point clouds from a scrambled Sobol sequence and then apply the coordinate-wise warping function $g(s)$ to create non-uniform distributions. The network is updated based on these generated, non-uniform clouds. This ensures that the model learns a general quadrature weight estimator adaptable to point density variations, rather than overfitting to a single point cloud.
The performance of the model is evaluated over a large number of independent trials ($M=2^{9}$). In each trial, a new, “unseen” point cloud is generated, and QuadrANN is used to compute the integral. The final results are the mean and standard deviation of these $M$ integral estimates. For a straightforward comparison, the standard Quasi-Monte Carlo (QMC) baseline computes its estimates on the corresponding $M$ original (unwarped) Sobol point clouds considering uniform weights. In all cases, the objective is to integrate a known function, allowing for explicit error analysis. The actual value of the integral is $1$ for the first and third examples, and $3/4$ for the second one.
5.1.1. Example 1: 2D Integration on a Unit Square
The first example serves as a reference for performance in a simple domain, $\Omega=[0,1]^{2}$. The integrand here is a 2D Gaussian probability density function (PDF) selected for its smoothness and localized characteristics:
$f_{1}(\mathbf{x})=\mathcal{N}(\mathbf{x}\,|\,\bm{\mu},\sigma^{2}\mathbf{I}).$ (13)
For this example, the parameters are set to $\bm{\mu}=(0.5,0.5)^{T}$ and $\sigma=0.025$ to center the function in the domain. The QuadrANN model was configured with $L=3$ layers, feature dimensions of $Q_{1}=2^{7}$, $Q_{2}=2^{8}$, $k=k^{\prime}=16$ nearest neighbors,positional encoding bands $p_{1}=3,p_{2}=5$, and $K_{\text{max}}=10$.
The results of this test case are summarized as follows. The QuadrANN model yielded a mean integral of $1.0004$ with a standard deviation of $0.0839$. These results compare to the Sobol–QMC method, which resulted in a mean of $0.9958$ and a standard deviation of $0.0964$. In the presented realization, QuadrANN achieved an absolute error of $0.0004$, while the corresponding Sobol–QMC error was $0.0042$. Moreover, our model demonstrated a higher precision by attaining about 12.912.9% reduction in standard deviation over the Sobol–QMC. Figure 1 visualizes the learned quadrature weights corresponding to a random “unseen” point cloud.
Figure 1. Interpolated quadrature weights learned by QuadrANN for the 2D unit square example. The color map indicates the weight value, and the dots refer to the locations of the points.
5.1.2. Example 2: 2D Integration on a Non-Convex Domain
To challenge the model with complex geometry, the second example is conducted on a non-convex, L-shaped domain, defined as $\Omega=[0,1]^{2}\setminus[0.5,1]^{2}$. The inner corner of this domain makes it particularly difficult for standard methods. The integrand is the same Gaussian PDF defined in Equation (13).
This setup tests the capacity of QuadrANN to adapt its weights to complex boundary. For this complex geometry, we increased the network depth to $L=4$. The additional propagation layer provides the model with a larger receptive field and a greater capacity for feature extraction, which is necessary to resolve the intricate non-convex boundary. The other model hyperparameters were set as in the previous case $Q_{1}=2^{7}$, $Q_{2}=2^{8}$, $k=k^{\prime}=16$, $p_{1}=3,p_{2}=5$, and $K_{\text{max}}=10$. Each point cloud consists of $n=2^{10}=1024$ points. It should be noted that in this test case, the base functions were re-normalized so that their integral over the L-shaped domain evaluates to one. The model is trained for $256$ epochs
For this test case, QuadrANN achieved a mean integral value of $0.7464$ with a standard deviation of $0.09164$. The Sobol–QMC produced a mean of $0.7434$ with a standard deviation of $0.09699$. Here, the QuadrANN model achieved an absolute error of $0.0036$, whereas the corresponding Sobol–QMC error was $0.0066$. Additionally, our model exhibited improved precision, achieving approximately a 5.55.5% reduction in standard deviation when compared to the Sobol–QMC approach.
The learned weights, shown in Figure 2, demonstrate the ability of the model to adjust its weights for adapting to the non-convex geometry.
Figure 2. Visualization of the quadrature weights for the L-shaped domain. The model accurately assigns the weight of the points to adapt to the non-convex geometry. The color map indicates the weight value, and the dots refer to the locations of the points.
5.1.3. Example 3: High-Dimensional Integration
The final pure integration example tests the performance of the model in a higher-dimensional space scenario, $\Omega=[0,1]^{4}$, where the curse of dimensionality is more pronounced. We compute the integral of a 4D Gaussian probability density function:
$f_{3}(\mathbf{x})=\mathcal{N}(\mathbf{x}\,|\,\bm{\mu},\sigma^{2}\mathbf{I}).$ (14)
with a mean $\bm{\mu}=(0.5,0.5,0.5,0.5)^{T}$, and a standard deviation $\sigma=0.075$. The network for this 4D case used $L=4$ layers, feature dimensions of $Q_{1}=2^{7}$, $Q_{2}=2^{8}$, $k=k^{\prime}=32$ neighbors, PE bands $p_{1}=5,p_{2}=7$, and $K_{\text{max}}=6$. The model is trained for $512$ epochs, using point clouds of size $n=2^{12}$.
In this high-dimensional test, QuadrANN achieved a mean integral of $0.9997$ with a standard deviation of $0.0993$. In contrast, standard Sobol-QMC produced a mean of $1.0020$ with a higher standard deviation of $0.1203$. In this example, QuadrANN achieved an absolute error of $0.0003$, while the associated Sobol–QMC error was $0.002$. Moreover, our model again exhibited enhanced accuracy by achieving approximately a 17.417.4% decrease in standard deviation compared to the Sobol–QMC.
5.1.4. Discussion on Stability
The results across all the examples underscore the benefits of our approach, particularly in reducing the variance of the integral estimator. The consistent reduction in the standard deviation of the integral estimates highlights a crucial feature of QuadrANN. While these examples address pure integration problems, the implications for extending to solve PDEs via functional minimization are clear. The optimization process at the core of such solvers seeks the minimum of a discretized integral. A quadrature rule with high variance provides a noisy estimate of this functional, which can mislead the optimization algorithm. By decreasing the variance, QuadrANN ensures a more stable loss behaviour for the optimizer. This stability is critical for ensuring the process converges robustly to the actual minimizer of the functional and yields a more accurate final solution.
5.2. Application to PDE Solvers
We now demonstrate the practical efficacy of our ANN-based quadrature rule within variational solvers for time-dependent PDEs. In each scenario, the task is to find the solution by minimizing a corresponding functional, where the accuracy of the numerical integration is critical. Through these applications, we aim to validate that a more accurate, data-driven quadrature rule leads to a more faithful optimization of the underlying functional and, consequently, a more precise solution. The variational time-stepping approach used here is inspired by minimizing movement schemes (IMEXs), which have proven effective in other complex deep learning contexts [19].
5.2.1. PDE Solution Network and Setup
For the forthcoming PDE examples, the time-dependent solution $u(t_{k},\mathbf{x})$ is approximated by a neural network, $U(t_{k},\mathbf{x};\theta)$. We employ a DGM-type architecture [20], detailed in Algorithm 2, which is proper for high-dimensional problems.
Algorithm 2 The DGM-type Solution Network Architecture
Input: Spatial coordinates $\mathbf{x}\in\mathbb{R}^{d}$, time $t$.
Learnable Parameters: $\{\mathbf{W}^{(\text{in})},\mathbf{b}^{(\text{in})},\{\mathbf{V}^{(s,l)},\mathbf{W}^{(s,l)},\mathbf{b}^{(s,l)}\}_{l=1}^{L},\mathbf{W}^{(\text{out})}\}$ for $s\in\{g,z,r,h\}$.
Hyperparameters: Number of layers $L$.
Output: Approximate solution $U(t,\mathbf{x};\theta)$.
procedure U($t,\mathbf{x}$)
  $\triangleright$ Input Layer
  $S^{(0)}\leftarrow\tanh(\mathbf{W}^{(\text{in})}[t,\mathbf{x}]^{T}+\mathbf{b}^{(\text{in})})$ $\triangleright$ Concatenate time and space
  $\triangleright$ DGM Hidden Layers
  for $l=1,\dots,L$ do
   $G^{(l)}\leftarrow\tanh(\mathbf{V}^{(g,l)}\mathbf{x}+\mathbf{W}^{(g,l)}S^{(l-1)}+\mathbf{b}^{(g,l)})$
   $Z^{(l)}\leftarrow\tanh(\mathbf{V}^{(z,l)}\mathbf{x}+\mathbf{W}^{(z,l)}S^{(l-1)}+\mathbf{b}^{(z,l)})$
   $R^{(l)}\leftarrow\tanh(\mathbf{V}^{(r,l)}\mathbf{x}+\mathbf{W}^{(r,l)}S^{(l-1)}+\mathbf{b}^{(r,l)})$
   $H^{(l)}\leftarrow\tanh(\mathbf{V}^{(h,l)}\mathbf{x}+\mathbf{W}^{(h,l)}(S^{(l-1)}\odot R^{(l)})+\mathbf{b}^{(h,l)})$
   $S^{(l)}\leftarrow(1-G^{(l)})\odot H^{(l)}+Z^{(l)}\odot S^{(l-1)}$ $\triangleright$ The new hidden state
  end for
  $\triangleright$ Estimated solution
  $U\leftarrow\text{Softplus}(\mathbf{W}^{(\text{out})}S^{(L)})$
  return $U$
end procedure
The solution network $U(t,\mathbf{x};\theta)$ is configured with $L=3$ hidden layers and a width of $32$ nodes per layer. For the 2D Heat Equation, we use a time step of $\tau=0.01$. The optimization for the first time step ($k=1$) is performed for $2^{11}$ epochs, while all subsequent steps are trained for $2^{7}$ epochs with a batch size of $8$. For the 4D Fokker-Planck equation, the time step is $\tau=0.01$. The first step is trained for $2^{11}$ epochs and the rest for $2^{7}$ epochs, using now larger point clouds of $n=2^{12}$ points, with a batch size of $4$.
5.2.2. Example: Heat Equation
We wish to solve the heat equation on $\Omega=[0,1]^{2}$. The problem is to find $u(t,\mathbf{x})$ for $t\in[0,1]$ and $\mathbf{x}=(x_{1},x_{2})\in\Omega$, such that:
$\begin{cases}\partial_{t}u-c\Delta u=f(t,\mathbf{x})&\text{in }(0,1]\times\Omega,\ c=1/8\\ u(0,\mathbf{x})=0&\text{in }\Omega\\ \frac{\partial u}{\partial\mathbf{n}}(t,\mathbf{x})=0&\text{on }(0,1]\times\partial\Omega\end{cases}$ (15)
The source term $f(t,\mathbf{x})$ is defined as:
$f(t,\mathbf{x})=\begin{cases}50\left(1+\cos\left(\frac{\pi r(\mathbf{x})}{R}\right)\right)\left(2+\cos(4\pi t)\right)e^{-3t}&\text{if }r(\mathbf{x})<R\\ 0&\text{if }r(\mathbf{x})\geq R\end{cases},$ (16)
where $r(\mathbf{x})=\sqrt{(x_{1}-0.5)^{2}+(x_{2}-0.5)^{2}}$ and $R=0.05$. We estimate the solution $u^{k}(\mathbf{x})$ at time $t_{k}=k\tau$ by minimizing the cost functional $\mathcal{J}[u^{k}]$:
$\mathcal{J}[u^{k}]=\int_{\Omega}\frac{1}{2}\left(u^{k}(x)-u^{k-1}(x)\right)^{2}+\frac{1}{2}\tau c\left(\nabla u^{k}(x)\right)^{2}-\tau f(t^{k},x)u^{k-1}(x)dx.$ (17)
The quadrature model is first trained on the $0,1]^{2}$ domain for $256$ epochs, where each batch consists of a point cloud of $2^{10}$ points created using the adopted warping policy. The trained model is then used to provide the quadrature weights for minimizing $\mathcal{J}[u^{k}]$. The solution obtained using QuadrANN integration achieved a mean absolute error of $1.527\times 10^{-3}$, a clear improvement over the error of $4.329\times 10^{-3}$ from the solver using standard Sobol-based Monte Carlo integration. This confirms the advantage of our approach for about 15% computational overhead. A qualitative comparison is presented in Figure 3.
Figure 3. Comparison of solutions for the 2D Heat Equation. A 1D slice of the computed solution $u(T,\mathbf{x})$ at the final time $T=1$, taken along the $x_{1}$-axis at $x_{2}=0.5$. The plot compares the solution obtained using our proposed QuadrANN integration with the standard Sobol-based Monte Carlo (MC) method against the Finite Difference Method (FDM) benchmark. The QuadrANN-based solution shows a closer agreement with the benchmark.
5.2.3. Example: Fokker-Planck Equation
To test our method in a higher-dimensional scenario, we focus on the Fokker-Planck equation within the $\Omega=[0,1]^{4}$.
The problem describes the probability density function, $u(t,\textbf{x})$, for the state of particle whose position $X_{t}$ evolves according to an Ornstein-Uhlenbeck type Stochastic Differential Equation (SDE):
$d\mathbf{X}_{t}=-\gamma(\mathbf{X}_{t}-\mathbf{x}_{0})dt+\sigma d\mathbf{W}_{t},$ (18)
where the drift is $\gamma=0.5$, the mean-reversion level is $\mathbf{x}_{0}=(0.5,0.5,0.5,0.5)^{T}$, the diffusion is $\sigma=0.2$, and $\mathbf{W}_{t}$ is a 4D Brownian motion. We assume again Neumann boundary conditions on $\partial\Omega$. The corresponding Fokker-Planck equation, which governs the time evolution of the probability density function $u(t,\mathbf{x})$, is given by:
$\partial_{t}u=\nabla\cdot[\gamma(\mathbf{x}-\mathbf{x}_{0})u]+\frac{\sigma^{2}}{2}\Delta u.$ (19)
The considered initial condition, $u(0,\mathbf{x})=u_{0}(\mathbf{x})$, is a mixture of two multivariate Gaussian distributions creating a bimodal landscape:
$u_{0}(\mathbf{x})=0.5\ \mathcal{N}(\mathbf{x}|\bm{\mu}_{1},\sigma_{1}^{2}\mathbf{I})+0.5\ \mathcal{N}(\mathbf{x}|\bm{\mu}_{2},\sigma_{2}^{2}\mathbf{I}),$ (20)
with means $\bm{\mu}_{1}=(0.35,0.35,0.35,0.35)^{T}$, $\bm{\mu}_{2}=(0.65,0.65,0.65,0.65)^{T}$, and standard deviations of $0.1$ and $0.11$, respectively.
Following a variational time-stepping scheme, we approximate the solution $u^{k}(\mathbf{x})$ at each time step by minimizing the cost functional $\mathcal{J}[u^{k}]$:
$\begin{split}\mathcal{J}[u^{k}]=\int_{\Omega}\bigg{[}&\frac{1}{2}\left(u^{k}(\mathbf{x})-u^{k-1}(\mathbf{x})\right)^{2}+\frac{\tau\sigma^{2}}{4}|\nabla u^{k}(\mathbf{x})|^{2}\\ &-\tau\gamma\left((\mathbf{x}-\mathbf{x}_{0})\cdot\nabla u^{k-1}(\mathbf{x})\right)u^{k}(\mathbf{x})-2\tau\gamma u^{k}(\mathbf{x})\bigg{]}d\mathbf{x}.\end{split}$ (21)
A benchmark solution is obtained by simulating the SDE for $10^{7}$ particles and using Kernel Density Estimation (KDE) with a bandwidth parameter of $0.03$, a standard non-parametric technique [21], on the final particle positions to construct a high-fidelity density. The error of our solver is quantified by the Hellinger distance to this benchmark.
For this 4D problem, a non-uniform point cloud is generated by applying a coordinate-wise warping function $g(s)$ to a 4D Sobol sequence. A dedicated QuadrANN model is trained for $512$ epochs on point clouds of $n=2^{12}$. This trained model then supplies the weights for the variational solver. The final solution using QuadrANN integration achieved a Hellinger distance of $0.0769$ to the KDE benchmark. This is considered an improvement over the Hellinger distance of $0.0917$ obtained with standard Sobol-QMC integration, achieved at a modest 20% increase in computational time. Figure 4 provides a visual comparison of the obtained solutions.
Figure 4. Comparison of solutions for the 4D Fokker-Planck equation. A 1D slice of the computed density $u(T,\mathbf{x})$ at the final time $T=1/2$, taken along the $x_{1}$-axis with other coordinates fixed at the center of the domain ($x_{2}=x_{3}=x_{4}=0.5$). The plot compares the density profile from our QuadrANN integration scheme with a standard Sobol-based Monte Carlo (Quasi-MC) method against the benchmark density constructed via SDE path simulation and Kernel Density Estimation (KDE).
6. Conclusions
In this work, we introduced QuadrANN, a deep learning architecture designed to overcome the critical challenge of numerical integration on non-uniform point clouds, a common task in modern mesh-free variational solvers for PDEs. Mainstream methods like Quasi-Monte Carlo struggle to handle such distributions, while the traditional permutation-variant neural networks make them unsuitable for unordered point sets. Our approach treats these limitations by adopting a Graph Neural Network (GNN) to learn data-driven, permutation-invariant quadrature rules directly from geometry. The QuadrANN architecture introduces a rich geometric message-passing scheme that combines absolute and relative positional features as well as a density feature in its initial layer. In contrast, the subsequent layers exploit a global context vector to make the learned rule aware of both local point density and the overall domain shape.
We demonstrated the efficiency of our method through a series of numerical experiments. Direct integration tests on simple, non-convex, and high-dimensional domains confirmed that QuadrANN consistently and significantly reduces the variance of the integral estimate compared to standard Sobol-QMC integration. The uncertainty reduction led to improved performance in solving PDE problems. In particular, when the quadrature model was integrated into variational ML-solvers for the Heat and Fokker-Planck equations, our method yielded more accurate final solutions than those obtained using standard QMC quadrature, validating its practical merit.
The core benefit of this work is that it provides a stable framework for the optimization process used in mesh-free variational solvers. Particularly, QuadrANN ensures a more reliable convergence to the energy minimizer by reducing the variance of the estimations of the integral, even while providing mean estimates of the integrals comparable to those from Sobol-QMC.
QuadrANN Implementation
We have developed QuadrANN using Pytorch-Geometric[22] on top of PyTorch[23]. Interested readers can refer to the following GitHub repository (github.com/kesmarag/QuadrANN) for the source code.